OpenStudy (anonymous):

solve using elimination 5x+3y=-13 7x-2y=19

OpenStudy (anonymous):

did u see how i got it I multiplied 1st equation by 2 and second by 3 to get common number at y

OpenStudy (anonymous):

x= 1

OpenStudy (anonymous):

dodomaroc, My answer needs to be in ordered pair??? It is saying it is wrong?? @hahd yes I have gotten that far, now pluging it into the first equation is where i need help

OpenStudy (anonymous):

y = -6

OpenStudy (anonymous):

thank yo dodomaroc !! :)

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

how did you figure out y???

OpenStudy (anonymous):

u can add the two equations: 12x + y = 6 y = 6 -12 =-6

OpenStudy (anonymous):

k